OpenStudy (anonymous):

Identify the open intervals on which the function is increasing or decreasing. (Select all that apply.) h(x) = 3x − x^3 Increasing: (−infinity, −1) (−1, 1) (1,+infinity ) none of these . Decreasing: (−infinty, −1) (−1, 1) (1,+infinity ) none of these . these r answer choices whichever apply should be ticked..so which ones should be

OpenStudy (anonymous):

How's is this different?

OpenStudy (anonymous):

jsut the wya i wrote it

OpenStudy (anonymous):

can u pleas ejust tell me which ones to select im getting timed..omg sorry

OpenStudy (anonymous):

please imran

OpenStudy (anonymous):

We need again to determine the derivative h'(x) = 3 - 3x^2 We'll detemrine the roots: 3-3x^2 = 0 1 - x^2 =0 (1-x)(1+x)=0 1-x=0 =>x=1 1+x=0 => x=-1 The function is decreasing over the intervals (-infinite,-1) and (1,infinite) and it is increasing over the interval [-1,1].